# When Experimental Economics Meets

### v1

@@ 5 Discussion
This paper establishes a methodological framework for LLM experiments by integrating principles from experimental economics and insights from artificial intelligence. Specifically, we propose eight actionable tactics for conducting LLM experiments. Our case studies with two sets of tasks and four models underscore the importance of these tactics. We find that assigned personas have a considerable impact on economic preferences but do not affect economic rationality. Furthermore, the single-turn dialogue type reduces the economic rationality of two open-source models with small parameter sizes, Llama and Qwen, but not for the others, GPT and DeepSeek. Similarly, discretizing the continuous choice set into multiple choices leads to a reduction in rationality for Llama and Qwen, as well as substantial changes in decisions in behavioral games for all LLMs. These highlight the sensitivity of LLMs’ performance to variations in experimental design and implementation.
Our paper contributes to the emerging literature at the intersection of LLMs and experimental economics, with a particular focus on advancing methodological approaches. Previous research mainly explores how LLMs can be effectively used as supplementary tools at various stages of human experiments [18], and how to use generative AI in experimental research [17]. In addition to these studies, we shift the perspective by treating LLMs as experimental objects and develop a framework for standardizing experimental protocols in LLM experiments. In addition, we extend the literature by presenting two case studies that demonstrate the application and validity of our proposed framework.
@@
Finally, we acknowledge several limitations and outline directions for future research emerging from this work. First, we focus exclusively on the individual decision making of LLMs. Extending this framework to accommodate multi-agent settings and human–AI hybrid experiments will be important for broadening its applicability. Second, our study proposes a set of protocols and evaluates them in the context of text-based interactions. As LLMs’ architectures continue to evolve, especially with the development of multimodal systems that integrate text, voice, image, and video data, further validation will be necessary to ensure the effectiveness of these protocols for next-generation models. Third, while we assess how design factors influence LLMs’ performance, the underlying mechanisms are not understood sufficiently. Enhancing the interpretability of LLMs’ decision-making is crucial to validate whether observed behavior reflects genuine human-like reasoning or simply mimics patterns in training data [70, 9, 8]. Addressing these issues will improve experimental protocols and observations, develop interpretations and theories of LLMs’ behavior, and support the effective application and integration of LLMs into everyday life.
@@
Table S1: The Description of Representative LLM Experiments
@@
Table S2: The Prompt Parameter Specification of Representative LLM Experiments
[b] Study Temperature Persona Example Understanding question Dialogue type Answer type Invalid answers Binz and Schulz [7] 0 Multi-turn Both Chen et al. [23] 0, 0.5, 1 Both Y Single-turn Both Hagendorff et al. [46] 0 -- Open Y Webb et al. [89] 0 Multi-turn Both Brookins and DeBacker [13] 1 Y -- Both Y Suri et al. [85] -- -- Both Goli and Singh [44] 1 -- Choice Macmillan-Scott and Musolesi [64] 1 -- Both Y Mei et al. [66] 1 Both Both Multi-turn Both Kosinski [55] 0, 1 -- Open Wang et al. [88] 0.7 Both -- Both Chen et al. [22] 1 -- Both Y Notes: In terms of temperature, we only summarize the temperature of models in GPT family for comparison, and “--” represents the parameter is not revealed in the study. “Y” represents the parameter or feature is specified in the study, while the black represents the opposite. “Both” represents that the study includes the parameter or feature in some experimental conditions or tasks and not in the others. For dialogue type, “--” represents the feature is not applicable, as each LLM agent only needs to answer one question. For answer type, “Open” represents the open-ended type and “Choice” represents the multiple-choice type. “Invalid answers” describes whether the study reports the proportion of invalid answers.
@@
Figure S1: Illustration of Dialogue Type
@@ A.1 Key Principles
To review key principles for human experiments, we resort to classical textbooks and survey papers that offer comprehensive discussions on experimental methodology. We choose two textbooks [30, 41], which have high citation counts. [^9] As of April 9, 2025, the Google Scholar citation counts for Davis and Holt [30] and Friedman and Sunder [41] were 3,197 and 1,442, respectively. Croson [28] is also included for its detailed comparison in experimental methods between economics and psychology. Two students who have taken the Ph.D. level experimental economics class independently read these three references and summarize the key methods for experimental design and implementation. We take the union of these principles and then distill them based on two criteria: 1) they are mentioned in at least two of the three sources, and 2) they have potential applicability to LLM experiments. Table S3 provides a summary of key guidelines from these three sources.
Table S3: Key Principals for Human Experiments
[b] DH1993 FS1994 Croson2005 Panel A: Basic rules in Experimental Design Follow induced value theory Avoid deception Minimize experimenter demand effects Panel B: Execution Rules in Experimental Implementation Use context-neutral instructions Make instructions comprehensible Provide illustrative examples Incorporate comprehension assessments Prevent inter-subject communication Notes: Croson2005 refers to Croson [28], DH1993 refers to Davis and Holt [30], and FS1994 refers to Friedman and Sunder [41]. A checkmark indicates that the corresponding rule is discussed in the reference.
@@
    Follow induced value theory. For economic experiments, a central consideration is the alignment of monetary incentives with induced value theory. Specifically, payment schemes must adhere to principles such as monotonicity, salience, and dominance [83]. Induced value theory ensures a linkage between participants’ payoff and their decisions, thereby eliciting truthful responses. In economic experiments, cash is typically used as the reward medium, instead of other methods, e.g., grades or extra credit points. [^10] Grades or extra credit points may not be valued by every student, whereas cash is usually non-satiable [28]. Moreover, though negative payment is possible in theory, subjects’ payments are always positive in practice.
@@ A.2 Sample Instructions
This is an experiment in the economics of market decision making. Various research foundations have provided funds for the conduct of this research. The instructions are simple, and if you follow them carefully and make good decisions you might earn a considerable amount of money which will be paid to you in cash after the experiment. [Make instructions comprehensible. Follow induced value theory.]
@@ B.1 A Brief History of LLMs
Figure S2: Timeline of Milestone in LLMs’ History: 2013 - 2023
Figure S2 presents a timeline of pivotal events in the evolution of large language models. The progression began with Word2Vec in the middle of 2013 [67], which established a foundation for learning semantic representations of words. The rapid evolution of LLMs can then be traced back to the introduction of the Transformer architecture by Vaswani et al. [86], which employed self-attention to overcome recurrent neural network (RNN) bottlenecks in processing long sequences. By parallelizing data handling instead of processing sequences step by step, Transformers significantly boosted efficiency and paved the way for subsequent breakthroughs.
@@ B.2 How to Type LLMs
In this section, we type LLMs by different features and discuss their implications for research. Table S4 presents the description of the representative models by types.
@@
Table S4: Main Features of Representative LLMs
@@ B.3 How to Prompt LLMs
The following sample code (Figure S3) demonstrates how to use the OpenAI API with specific settings to generate a response to a decision-making task. In an interaction with API calls or web-based chat interfaces, there are generally two key aspects that users need to adjust. The first is the specification of hyperparameters, which govern the model’s output behavior. The second aspect involves the formulation of the prompt itself, which serves as the foundation of the interaction. Proper tuning of both the hyperparameters and the prompt is essential to obtain meaningful and reliable output. This section will explore these two critical components in detail.
Figure S3: Sample Code for Prompting LLMs
@@
Randomness in outputs: Sampling-based decoding and temperature. The temperature parameter is central to controlling the randomness of a LLM’s output. Technically, temperature scales the model’s logits—raw scores computed for each potential token—prior to applying the softmax function. [^13] Specifically, $p_{i}=\frac{\exp\left(\frac{z_{i}}{T}\right)}{\sum_{j}^{n}\exp\left(\frac{z_{j}}{T}\right)}$, where $p_{i}$ denotes the probability of choosing the token, $z_{i}$ represents the logit of token $i$ and $T$ is the temperature. When $T=0$, $p_{i}=1$ if $z_{i}=\max\{z_{1},\cdots,z_{n}\}$ and $p_{i}=0$ if $z_{i}\neq\max\{z_{1},\cdots,z_{n}\}$. When $T\rightarrow\infty$, $p_{i}\rightarrow\frac{1}{n}$. Lower temperature values sharpen the probability distribution, making high-probability tokens far more likely to be selected. This results in outputs that are more deterministic, coherent, and consistent. In contrast, higher temperatures flatten the distribution, increasing the chances of selecting less likely tokens and promoting more diverse, or unexpected outputs. In practice, temperature selection involves a fundamental trade-off between predictability and exploration. For tasks requiring strict adherence to factual or logical structure (e.g., legal document generation), low temperatures minimize variability and reduce error rates. Conversely, higher temperatures are often necessary for open-ended tasks where diverse outputs are desirable, such as generating multiple hypotheses in exploratory research.
@@ Appendix C Additional Analysis in Case Study 1
Temperature   We examine the temperature from 0 to 1 with a step of 0.1, and plot the differences in average CCEI scores between each condition and the baseline with the temperature of 1 in Figure S4. We do not observe any significant effect of temperature on CCEI scores in both risk and social preference domains (all $p>0.1$). Our findings are aligned with the literature and further support Tactic 1.
Incentive   As induced value theory emphasizes the crucial role of monetary incentives in human experiments, we assess whether the existence of hypothetical incentives and stake sizes affect economic rationality in LLM experiments. We perform the following conditions independently. First, we remove the payment scheme from the prompts. Second, we enlarge the stakes by 10, 100 and 1,000 times separately to examine the effect of stake sizes. Section D.2.2 details the prompts and Figure S4 presents the results.
@@
Figure S4: Comparison in rationality of LLMs. The baseline condition is omitted in the figure, where the corresponding temperature is 1 for GPT, 0.3 for DeepSeek, 0.6 for Llama and 0.7 for Qwen. The markers represent differences in average CCEI scores between experimental and baseline conditions. The error bars represent 95% confidence intervals.
Example   We evaluate the impact of the example by removing it from the baseline condition. As illustrated in Figure S4, no significant differences are observed in the risk preference domain. In terms of social preference, GPT generates slightly higher CCEI when removing the example, while DeepSeek and Qwen result in lower CCEI scores compared to the baseline (all $p<0.01$).

### v2

@@ 6 Discussion
Integrating established principles from experimental economics with insights from artificial intelligence, this paper systematically investigates how key considerations from experimental design to implementation influence LLMs’ outputs. We find that assigned personas have a considerable impact on economic preferences but do not affect economic rationality. Furthermore, the single-turn dialogue type reduces the economic rationality of two open-source models with small parameter sizes, Llama and Qwen, but not for the others, GPT and DeepSeek. Similarly, discretizing the continuous choice set into multiple choices leads to a reduction in rationality for Llama and Qwen, as well as substantial changes in decisions in behavioral games for all LLMs. These highlight the sensitivity of LLMs’ performance to variations in experimental design and implementation. Based on these observations, we summarize seven tactics for designing and conducting LLM experiments, thereby contributing to more comparable and replicable research practices.
Our paper adds to the emerging literature at the intersection of LLMs and experimental economics, with a particular focus on advancing methodological approaches. Previous research mainly explores how LLMs can be effectively used as supplementary tools at various stages of human experiments [18], and how to use generative AI in experimental research [17]. In addition to these studies, we shift the perspective by treating LLMs as experimental objects and develop a framework for standardizing experimental protocols in LLM experiments. Moreover, this framework is grounded in two case studies, which further enhances its validity.
@@
Finally, we acknowledge several limitations and outline directions for future research emerging from this work. First, we focus exclusively on the individual decision making of LLMs. Extending this framework to accommodate multi-agent settings and human–AI hybrid experiments [36] will be important for broadening its applicability. Second, our study proposes a set of protocols and evaluates them in the context of text-based interactions. As LLMs’ architectures continue to evolve, especially with the development of multimodal systems that integrate text, voice, image, and video data, further validation will be necessary to ensure the effectiveness of these protocols for next-generation models. Third, while we assess how design factors influence LLMs’ performance, the underlying mechanisms are not understood sufficiently. Enhancing the interpretability of LLMs’ decision-making is crucial to validate whether observed behavior reflects genuine human-like reasoning or simply mimics patterns in training data [73, 9, 8]. Addressing these issues will improve experimental protocols and observations, develop interpretations and theories of LLMs’ behavior, and support the effective application and integration of LLMs into everyday life.
@@
Supplementary Table 1: The Description of Representative LLM Experiments
@@
Supplementary Table 2: The Prompt Parameter Specification of Representative LLM Experiments
[b] Study Temperature Persona Incentive Example Understanding question Dialogue type Answer type Invalid answers Binz and Schulz [7] 0 Both Multi-turn Both Chen et al. [23] 0, 0.5, 1 Both Y Y Single-turn Both Hagendorff et al. [47] 0 -- Open Y Webb et al. [92] 0 Multi-turn Both Brookins and DeBacker [13] 1 Y Y -- Both Y Suri et al. [88] -- -- Both Goli and Singh [45] 1 Y -- Choice Macmillan-Scott and Musolesi [66] 1 -- Both Y Mei et al. [68] 1 Both Y Both Multi-turn Both Kosinski [57] 0, 1 -- Open Wang et al. [91] 0.7 Both -- Both Chen et al. [22] 1 Both -- Both Y Notes: In terms of temperature, we only summarize the temperature of models in GPT family for comparison, and “--” represents the parameter is not revealed in the study. “Y” represents the parameter or feature is specified in the study, while the black represents the opposite. “Both” represents that the study includes the parameter or feature in some experimental conditions or tasks and not in the others. For dialogue type, “--” represents the feature is not applicable, as each LLM agent only needs to answer one question. For answer type, “Open” represents the open-ended type and “Choice” represents the multiple-choice type. “Invalid answers” describes whether the study reports the proportion of invalid answers.
@@
Supplementary Figure 1: Illustration of Dialogue Type
@@ A.1 Key Principles
To review key principles for human experiments, we resort to classical textbooks and survey papers that offer comprehensive discussions on experimental methodology. We choose two textbooks [30, 42], which have high citation counts. [^9] As of April 9, 2025, the Google Scholar citation counts for Davis and Holt [30] and Friedman and Sunder [42] were 3,197 and 1,442, respectively. Croson [28] is also included for its detailed comparison in experimental methods between economics and psychology. Two students who have taken the Ph.D. level experimental economics class independently read these three references and summarize the key methods for experimental design and implementation. We take the union of these principles and then distill them based on two criteria: 1) they are mentioned in at least two of the three sources, and 2) they have potential applicability to LLM experiments. Table 3 provides a summary of key guidelines from these three sources.
Supplementary Table 3: Key Principals for Human Experiments
DH1993 [30] FS1994 [42] Croson2005 [28]
Panel A: Basic Rules in Experimental Design
Follow the induced value theory
Avoid deception
Minimize experimenter demand effects
Panel B: Execution Rules in Experimental Implementation
Use context-neutral instructions
Make instructions comprehensible
Provide illustrative examples
Incorporate comprehension assessments
Prevent inter-subject communication
@@
    Follow the induced value theory. For economic experiments, a central consideration is the alignment of monetary incentives with induced value theory. Specifically, payment schemes must adhere to principles such as monotonicity, salience, and dominance [86]. Induced value theory ensures a linkage between participants’ payoff and their decisions, thereby eliciting truthful responses. In economic experiments, cash is typically used as the reward medium, instead of other methods, e.g., grades or extra credit points. [^10] Grades or extra credit points may not be valued by every student, whereas cash is usually non-satiable [28]. Moreover, though negative payment is possible in theory, subjects’ payments are always positive in practice.
@@ A.2 Sample Instructions
This is an experiment in the economics of market decision making. Various research foundations have provided funds for the conduct of this research. The instructions are simple, and if you follow them carefully and make good decisions you might earn a considerable amount of money which will be paid to you in cash after the experiment. [Make instructions comprehensible. Follow the induced value theory.]
@@ B.1 A Brief History of LLMs
Supplementary Figure 2: Timeline of Milestones in LLMs’ History: 2013 - 2023
Figure 2 presents a timeline of pivotal events in the evolution of large language models. The progression began with Word2Vec in the middle of 2013 [70], which established a foundation for learning semantic representations of words. The rapid evolution of LLMs can then be traced back to the introduction of the Transformer architecture [89], which employed self-attention to overcome recurrent neural network (RNN) bottlenecks in processing long sequences. By parallelizing data handling instead of processing sequences step by step, Transformers significantly boosted efficiency and paved the way for subsequent breakthroughs.
@@ B.2 How to Type LLMs
In this section, we type LLMs by different features and discuss their implications for research. Table 4 presents the description of the representative models by types.
@@
Supplementary Table 4: Main Features of Representative LLMs
@@ B.3 How to Prompt LLMs
In an interaction with LLMs using API calls or web-based chat interfaces, there are generally two key aspects that users need to adjust. The first is the specification of hyperparameters, which govern the model’s output behavior. The second aspect involves the formulation of the prompt itself, which serves as the foundation of the interaction. Proper tuning of both the hyperparameters and the prompt is essential to obtain meaningful and reliable output. This section will explore these two critical components in detail.
@@
Randomness in outputs: Sampling-based decoding and temperature. The temperature parameter is central to controlling the randomness of LLMs’ output. Technically, temperature scales the model’s logits—raw scores computed for each potential token—prior to applying the softmax function. [^13] Specifically, $p_{i}=\frac{\exp\left(\frac{z_{i}}{T}\right)}{\sum_{j}^{n}\exp\left(\frac{z_{j}}{T}\right)}$, where $p_{i}$ denotes the probability of choosing the token, $z_{i}$ represents the logit of token $i$ and $T$ is the temperature. When $T=0$, $p_{i}=1$ if $z_{i}=\max\{z_{1},\cdots,z_{n}\}$ and $p_{i}=0$ if $z_{i}\neq\max\{z_{1},\cdots,z_{n}\}$. When $T\rightarrow\infty$, $p_{i}\rightarrow\frac{1}{n}$. Lower temperature values sharpen the probability distribution, making high-probability tokens far more likely to be selected. This results in outputs that are more deterministic, coherent, and consistent. In contrast, higher temperatures flatten the distribution, increasing the chances of selecting less likely tokens and promoting more diverse, or unexpected outputs. In practice, temperature selection involves a fundamental trade-off between predictability and exploration. For tasks requiring strict adherence to factual or logical structure (e.g., legal document generation), low temperatures minimize variability and reduce error rates. Conversely, higher temperatures are often necessary for open-ended tasks where diverse outputs are desirable, such as generating multiple hypotheses in exploratory research.
@@ Appendix C Additional Analysis in Case Study 1
Temperature   We examine the temperature from 0 to 1 with a step of 0.1, and plot the differences in average CCEI scores between each condition and the baseline with the temperature of 1 in Figure 3. We do not observe any significant effect of temperature on CCEI scores in both risk and social preference domains (all $p>0.1$).
Incentive   As induced value theory emphasizes the crucial role of monetary incentives in human experiments, we assess whether the existence of hypothetical incentives and stake sizes affect economic rationality in LLM experiments. We perform the following conditions independently. First, we remove the payment scheme from the prompts. Second, we enlarge the stakes by 10, 100 and 1,000 times separately to examine the effect of stake sizes. Section D.2.2 details the prompts and Figure 3 presents the results.
@@
Supplementary Figure 3: Comparison in rationality of LLMs. The baseline condition is omitted in the figure, where the corresponding temperature is 1 for GPT, 0.3 for DeepSeek, 0.6 for Llama and 0.7 for Qwen. The markers represent differences in average CCEI scores between experimental and baseline conditions. The error bars represent 95% confidence intervals.
Example   We evaluate the impact of the example by removing it from the baseline condition. As illustrated in Figure 3, no significant differences are observed in the risk preference domain. In terms of social preference, GPT generates slightly higher CCEI when removing the example, while DeepSeek and Qwen result in lower CCEI scores compared to the baseline (all $p<0.01$).
@@ D.1.1 Baseline Condition
In subsequent rounds after the first round of each simulation, the system prompts remain the same, and the input incorporates both the prior user message and LLMs’ responses (Figure 1(b)). The new question in the user message for the current round is listed below.